OpenStudy (anonymous):

1/3x +1/3y = 3 Which of the following equations is equivalent to the given equation? 3x + 2y = 3 2x + 3y = 18 3x + 2y = 18

OpenStudy (anonymous):

HELP PLEASE!

OpenStudy (anonymous):

\[\frac{ 1 }{ 3 }x+\frac{ 1 }{ 3 }y =3\] multiply by 3 and get x + y =9 so either I misunderstood the question, you typed it wrong, or there is another choice